Determine the corresponding t values for the indicated area assume df=5. NOT TO SCALEDetermine the corresponding t values for the indicated area assume df -5. NOT TO SCALE SELECT ALL APPLICAELE CHOICES A) B) t 2.570582 t 2.641 0.3 df-5 a area -0.025 C) D) None of These 0.2 t 2.631 0.1
Answer #1

Given that,

Degrees of freedom = DF = 5

Significance level (\alpha) = 0.025

We want to find, t value for area = 0.025 with df = 5

Using Excel we get, t value = 2.570582

Excel command:

TINV(2 * 0.025, 5) = TINV(0.05, 5) = 2.570582

Answer: A) t = 2.570582